OpenStudy (anonymous):

a + 5 1 3a - 1 ---- - -- = ------- 3 2 4

OpenStudy (anonymous):

get common denominator - 12 (in this case) so, multiply both sides by 12, so 4*(a+5)-6=3*(3a-1) 4a+20-6=9a-3 can you finish it?